Interference waves of sound

1. Two loudspeakers 6.0 m apart are playing the same frequency. If you stand 10.0 m in front of the plane of the speakers, centered between them, you hear a sound ofmaximum intensity. As you walk parallel to the plane of the speakers, staying 10.0 in front of them, you first hear a minimum of sound intensity when you are directlyin front of one of the speakers. What is the frequency of the sound? Assume a sound speed of 340 m/s.

2. Two in-phase loudspeakers emit identical 1000 Hz sound waves along the x-axis. What distance should one speaker be placed behind the other for the sound to have anamplitude 1.10 times that of each speaker alone?

3. Two radio antennas are separated by 1.60 m . Both broadcast identical 750 MHz waves. If you walk around the antennas in a circle of radius 20.0 m , how many maximawill you detect?
Answer #1
The distance between two loudspeakers = 6 mThe distance of person who stand infront of the speakers = 10 mThen the path of two waves = sqrt(6^2 + 10^2) = 11.7 m and 10 mSo the pathdifference between the waves = 11.7 - 10 = 1.7 mSo the wavelength = 1.7 * 2 = 3.4 mSpeed of sound = 340 m/sThen the frequency = v/λ = 340 / 3.4 = 100 Hz
answered by: Shaira